# Experimental Sensitivity Enhancement of a Quantum Rydberg Atom-Based RF Receiver with a Metamaterial GRIN Lens

### v1

@@ Abstract
We experimentally demonstrate enhanced sensitivity of an atom-based Rydberg radio frequency (RF) receiver integrated with a gradient refractive index (GRIN) Luneburg-type metamaterial lens. By analyzing the electromagnetically induced transparency (EIT) effect in Cesium vapor, we compare receiver performance with and without the GRIN lens under a 2.2 GHz and a 3.6 GHz far-field excitation. Our measurements reveal a significant amplification of the EIT transparency window when the lens is introduced, consistent with the theoretical prediction that the local E-field enhancement at the vapor cell reduces the minimum detectable electric field and increases the signal-to-noise ratio (SNR) of the Rydberg RF receiver. This experimental validation highlights the potential of metamaterial-assisted quantum sensing to overcome the inherent bandwidth and sensitivity limitations of bare Rydberg receivers for a variety of applications, such as electromagnetic compatibility (EMC) testing, quantum radar, and wireless communications.
@@
Atom-based radio frequency (RF) sensing utilizes the effect of electromagnetically induced transparency (EIT) in a non-linear quantum optical effect bHarris. EIT is used to excite the outermost electron, typically in a group 1 Alkali metal (such as Rubidium, Rb, or Cesium, Cs, which was used in this work), into a Rydberg state via a two-photon optical transition. Once the atom is excited to a Rydberg state, its unique optical and atomic properties enable the analysis of various phenomena, including studies of atomic structure and highly sensitive detection of electromagnetic waves bCai2022; bHolloway2014, terahertz imaging bKevin2016, quantum memory bKuzmich, slowed light bSlowlight, quantum computing bGiraldo, electric field metrology bEMC, quantum radar bRadar, and wireless communication bMenchetti; bOtto2021.
@@
While the theory of atom-based RF sensing promises extremely wide bandwidths, Planck constant level of precision, and a sub-noise floor signal detection, its real-world performance is hindered by the Doppler effect between multi-body Rydberg atoms, that is always present in hot atomic vapor cells (i.e., room-temperature practical systems), resulting in the receiver’s sensitivity being far from its theoretical expectation bSensitivity. Several methods have been considered in the literature to improve the sensitivity and the signal-to-noise ratio (SNR) of atom-based RF receivers, such as the use of a split ring resonator (SRR) bSRR to capture the incident RF field and enhance it at the resonant frequency of the SRR, or other types of resonant structures bSRR2; bSRR3. Despite being very effective, these methods cannot be considered suitable for electric field metrology due to inevitable spurious and harmonic emissions caused by resonating metals. Additionally, the narrowband nature of the resonating frequency further limits its potential applications. Other techniques consider the application of a Mach-Zehnder interferometer bLasers1 and combined laser arrays with cascaded diffraction gratings bLasers2. However, such techniques are complex and expensive to implement, while their gains are also confined to a narrow bandwidth.
@@
Therefore, we propose a fundamentally different approach for enhancing the sensitivity of a Rydberg RF receiver in this work, which has not been explored in the literature to the best of our knowledge. We propose encapsulating its vapor cell with a passive gradient refractive index (GRIN) metamaterial lens that does not consume active power nor create any spurious emissions. Since the receiver is expected to be in the far-field of the RF emission or wireless communication signal source, we consider a Luneburg-type GRIN lens that focuses planar wavefronts on a focal point, assumed to be at the center of the vapor cell, with a high gain and an ultrawide bandwidth. We designed, fabricated, and characterized the lens in an anechoic chamber. Then, we performed experimental measurement of the EIT window with a Rydberg RF receiver, reporting a significant improvement in its sensitivity.
@@
For a two-state Rydberg RF receiver, the minimal detectable electric field corresponding to the quantum shot noise in the Autler–Townes limit bBussey; bShot2015 can be expressed as:
@@
In this work, we consider a Luneburg-type gradient index (GRIN) lens with refractive index variation $n(r)=\sqrt{\left(2-\left(r/R\right)^{2}\right)}$ in a general radial coordinate $(r)$ that has the electric field focusing ability expressed analytically per bMaci:
@@
When an RF wave passes through the GRIN lens, its phase is adjusted to focus at the vapor cell center, producing a beam-focusing gain (Fig. 2). Initially, we measured the fabricated lens’s focusing gain $\left(\gamma\right)$ and its beam waist at 3.6 GHz in an anechoic chamber. Then, we tested the lens within the Rydberg atom-based RF receiver setup, comparing EIT splitting with and without the lens at 2.2 GHz and 3.6 GHz. To characterize the lens inside an anechoic chamber, we used an ETS-Lindgren model 3102 conical log spiral antenna with 3.5 dBi gain at 3.6 GHz, engaging both polarizations simultaneously. The antenna was placed at a 3.2 m distance to ensure that the far-field conditions ar met. Then, a coaxial line near-field probe was used to capture the E-field at various positions on the surface in a symmetrical x and y axes matrix, and up to 50 mm distance away in the z-axis. Results are shown in Fig. 3. We measured $\gamma$ of up to 8.42 dB at the focal point of the lens, gradually decreasing in x,y, and z directions away from it, in line with the theoretical diffraction limit bEhsan. The atom-based Rydberg RF receiver, described in this work and shown in Fig. 2, consists of:
@@
In this setup, the RF detection is performed with a balanced photodiode. The probe beam is then split by using a calcite beam displacer, where one beam is then counter-propagated with the coupling laser, and the other one traverses the cell as a reference. Aiming to emulate a typical linearly-polarized E-field source, we separately generated a 2.2 GHz and a 3.6 GHz unmodulated signal with 11 dBm transmit power and the Astro Antenna Ltd. AHA-118S double ridged horn antenna with $\approx$ 2 dBi and 5 dBi gain at these frequencies. The antenna was placed at a 2.2 m distance, and the lens was placed 24 mm away from the vapor cell’s center. The antenna was tilted 45 to create a slant angle, allowing the lens to enhance the RF signal in both polarizations. The experimental measurement result is shown in Fig. 4. It can be seen that the amount of EIT splitting has effectively doubled with the lens at both 2.2 GHz and 3.6 GHz, therefore significantly increasing the SNR of the Rydberg RF receiver across an ultrawide bandwidth.
In this letter, we have considered the application of metamaterial GRIN lens technology to enhance the sensitivity of atom-based Rydberg RF receivers without compromising the measurement accuracy by RF parasitics, such as spurious or harmonic emissions. We 3D printed a Luneburg-type GRIN metamaterial lens and demonstrated its effectiveness at increasing the SNR of the Rydberg RF receiver over an ultrawide bandwidth. The proposed solution is low-cost and suitable for various applications, including electric-field metrology, quantum radar, and wireless communications.

### v2

@@ Abstract
We experimentally demonstrate enhanced sensitivity of an atom-based Rydberg radio frequency (RF) receiver integrated with a gradient refractive index (GRIN) Luneburg-type metamaterial lens. By analyzing the electromagnetically induced transparency (EIT) effect in Cesium vapor, we compare receiver performance with and without the GRIN lens under a 2.2 GHz and a 3.6 GHz far-field excitation. Our measurements reveal a significant amplification of the EIT window when the lens is introduced, consistent with the theoretical prediction that the local E-field enhancement at the vapor cell reduces the minimum detectable electric field and improves the microwave electric field measurement sensitivity of the Rydberg atom-based RF receiver over an ultrawide bandwidth of the lens. This experimental validation demonstrates the potential of metamaterial-enhanced quantum RF sensing for a wide range of applications, such as electromagnetic compatibility (EMC) testing, quantum radar, and wireless communications.
@@
Atom-based radio frequency (RF) sensing utilizes the effect of electromagnetically induced transparency (EIT) in a non-linear quantum optical effect Harris, S. E., et al. (1990). EIT is used to excite the outermost electron, typically in a group 1 Alkali metal (such as Rubidium, Rb, or Cesium, Cs, which was used in this work), into a Rydberg state via a two-photon optical transition. Once the atom is excited to a Rydberg state, its unique optical and atomic properties enable the analysis of various phenomena, including studies of atomic structure and highly sensitive detection of electromagnetic waves M. Cai, et al. (2022); C. L. Holloway, et al. (2014), terahertz imaging C. G. Wade, et al. (2017), quantum memory L. Li and A. Kuzmich (2016), slowed light R. Liu, et al. (2017), quantum computing F. Giraldo, et al. (2022), electric field metrology C. L. Holloway, et al. (2017), quantum radar S. Borówka, et al. (2024), and wireless communication M. Menchetti, et al. (2023); J. S. Otto, et al. (2021).
@@
While the theory of atom-based RF sensing promises extremely wide bandwidths, Planck constant level of precision, and a sub-noise floor signal detection, its real-world performance is hindered by the Doppler effect between multi-body Rydberg atoms, that is always present in hot atomic vapor cells (i.e., room-temperature practical systems), resulting in the receiver’s sensitivity being far from its theoretical expectation N. Prajapati, et al. (2023). Several methods have been considered in the literature to improve the sensitivity of bare Rydberg atom-based RF receivers, such as the use of a split ring resonator (SRR) C. L. Holloway, et al. (2022) to capture the incident RF field and enhance it at the resonant frequency of the SRR, or other types of resonant structures G. Sandidge, et al. (2024); B. Wu, et al. (2025a). Despite being very effective, these methods cannot be considered suitable for electric field metrology due to inevitable spurious and harmonic emissions caused by resonating metals. Additionally, the narrowband nature of the resonating frequency further limits its potential applications. Other techniques consider the application of a Mach-Zehnder interferometer W. Yang, et al. (2023) and combined laser arrays with cascaded diffraction gratings B. Wu, et al. (2025b). However, such techniques are complex and expensive to implement, while their gains are also confined to a narrow RF bandwidth.
@@
Therefore, we propose a fundamentally different approach for enhancing the sensitivity of a Rydberg RF receiver in this work, which has not been explored in the literature to the best of our knowledge. We propose encapsulating its vapor cell with a passive gradient refractive index (GRIN) metamaterial lens that does not consume active power nor create any spurious emissions. Since the receiver is expected to be in the far-field of the RF emission or wireless communication signal source, we consider a Luneburg-type GRIN lens that focuses planar wavefronts on a focal point, assumed to be at the center of the vapor cell, with a high gain and an ultrawide RF bandwidth. In this design, graded index profile arises via engineered subwavelength structure rather than continuous homogeneous optics. We designed, fabricated, and characterized the lens in an anechoic chamber. Then, we performed experimental measurement of the EIT window with a Rydberg RF receiver, reporting a significant improvement in its sensitivity.
@@
For a two-state Rydberg RF receiver, the minimal detectable electric field corresponding to the quantum shot noise in the Autler–Townes limit L. W. Bussey, et al. (2022); H. Fan, et al. (2015) can be expressed as:
@@
In this work, we consider a Luneburg-type gradient index (GRIN) lens with refractive index variation $n(r)=\sqrt{\left(2-\left(r/R\right)^{2}\right)}$ in a general radial coordinate $(r)$ that has the electric field focusing ability expressed analytically per I. Gashi, et al. (2024):
@@
When an RF wave passes through the GRIN lens, its phase is adjusted to focus at the vapor cell center, producing a beam-focusing gain (Fig. 2). Initially, we measured the fabricated lens’s focusing gain $\left(\gamma\right)$ and its beam waist at 3.6 GHz in an anechoic chamber. Then, we tested the lens within the Rydberg atom-based RF receiver setup, comparing EIT splitting with and without the lens at 2.2 GHz and 3.6 GHz. To characterize the lens inside an anechoic chamber, we used an ETS-Lindgren model 3102 conical log spiral antenna with 3.5 dBi gain at 3.6 GHz, engaging both polarizations simultaneously. The antenna was placed at a 3.2 m distance to ensure that the far-field conditions ar met. Then, a coaxial line near-field probe was used to capture the E-field at various positions on the surface in a symmetrical x and y axes matrix, and up to 50 mm distance away in the z-axis. Results are shown in Fig. 3. We measured $\gamma$ of up to 8.42 dB at the focal point of the lens, gradually decreasing in x,y, and z directions away from it, in line with the theoretical diffraction limit Hosseininejad, S.E., et al. (2019). The atom-based Rydberg RF receiver, described in this work and shown in Fig. 2, consists of:
@@
In this setup, the RF detection is performed with a balanced photodiode. The probe beam is then split by using a calcite beam displacer, where one beam is then counter-propagated with the coupling laser, and the other one traverses the cell as a reference. Aiming to emulate a typical linearly-polarized E-field source, we separately generated a 2.2 GHz and a 3.6 GHz unmodulated signal with 11 dBm transmit power and the Astro Antenna Ltd. AHA-118S double ridged horn antenna with $\approx$ 2 dBi and 5 dBi gain at these frequencies. The antenna was placed at a 2.2 m distance, and the lens was placed 24 mm away from the vapor cell’s center. The antenna was tilted 45 to create a slant angle, allowing the lens to enhance the RF signal in both polarizations. The experimental measurement result is shown in Fig. 4. It can be seen that the amount of EIT splitting has effectively doubled with the lens at both 2.2 GHz and 3.6 GHz, therefore increasing the sensitivity of the Rydberg RF receiver over an ultrawide bandwidth of the lens.
In this letter, we have considered the application of metamaterial GRIN lens technology to enhance the sensitivity of atom-based Rydberg RF receivers without compromising the measurement accuracy by RF parasitics, such as spurious or harmonic emissions. We 3D printed a Luneburg-type GRIN metamaterial lens and demonstrated its effectiveness at increasing the sensitivity of the Rydberg RF receiver over an ultrawide bandwidth of the lens. The proposed solution is low-cost and suitable for various applications, including electric-field metrology, quantum radar, and wireless communications.